OpenStudy (anonymous):

find the equation of a circle whose centre is (-3,-2) radius is 4

OpenStudy (anonymous):

should be \[\left( x--3 \right)^{2}+\left( y--2 \right)^{2}=4^{2}\] if i am not mistaken you should get \[x ^{2}+y ^{2}+6x+4y=3\]

OpenStudy (anonymous):

\[x^2-6x+y^2+4y=3\] check the signs and thanx